I have a question about midpoints and segment congruence. Here it is:
It says given the coordinate of one endpoint line AB and its midpoint M, find the coordinates of the other end point. A(0,9), M(2,5)

Answer by checkley71(8403)

a to m = 0+2=2 (x value)
a to m =9-5=4 (y value)
to find the end point we have to add the x distance to 2+2=4 & -4 to the y point 5-4=1 so the end point is (4,1)